OpenStudy (anonymous):

The graph of a parabola is represented by the equation y=ax^2 where a is a positive integer. If a is multiplied by 2, the new parabola would become? 1. Narrower and open downward 2. Narrower and open upward 3.wider and open downward 4. Open and open upward

OpenStudy (anonymous):

2

OpenStudy (anonymous):

How?

OpenStudy (campbell_st):

do a simply table of values a =1 so its y = x^2 x: -2 : -1 : 0 : 1 : 2 ---------------------- y: 4 : 1 : 0 : 1 : 4 now let a = 2, so its y = 2x^2 x: -2 : -1 : 0 : 1 : 2 ---------------------- y: 8 : 2 : 0 : 2 : 8 plot the points and the 2nd parabola is drawn inside the 1st parabola... with a common vertex of (0,0) the basic rule is for y = ax^2 a determines concavity in the parabola as well as the larger the value of a the steeper the graph of the parabola so consider y = 1/2 x^2 x: -2 : -1 : 0 : 1 : 2 ---------------------- y: 2 : 1/2 : 0 : 1/2 : 2 the is outside the 1 above.

OpenStudy (anonymous):

"the" What exactly is classified as "the"? (located at the last line)

OpenStudy (anonymous):

So the answer is basically choice #2?